# Learning deep illumination-robust features from multispectral filter array images

### v2

@@ Abstract
Multispectral (MS) snapshot cameras equipped with a MS filter array (MSFA), capture multiple spectral bands in a single shot, resulting in a raw mosaic image where each pixel holds only one channel value. The fully-defined MS image is estimated from the raw one through demosaicing, which inevitably introduces spatio-spectral artifacts. Moreover, training on fully-defined MS images can be computationally intensive, particularly with deep neural networks (DNNs), and may result in features lacking discrimination power due to suboptimal learning of spatio-spectral interactions. Furthermore, outdoor MS image acquisition occurs under varying lighting conditions, leading to illumination-dependent features. This paper presents an original approach to learn discriminant and illumination-robust features directly from raw images. It involves: raw spectral constancy to mitigate the impact of illumination, MSFA-preserving transformations suited for raw image augmentation to train DNNs on diverse raw textures, and raw-mixing to capture discriminant spatio-spectral interactions in raw images. Experiments on MS image classification show that our approach outperforms both handcrafted and recent deep learning-based methods, while also requiring significantly less computational effort.
Index Terms:
Multispectral imaging, MSFA, texture features, deep learning
I Introduction
Much of the recent success in several pattern recognition applications, such as crop/weed recognition [3, 33] and texture classification [28, 2], is thanks in part to the advancement of cameras and specifically to the emergence of compact ”snapshot” MS cameras [16]. These cameras rely on an MSFA laid over the sensor to acquire several spectral bands in one shot and provide a raw mosaic image, in which a single channel value is available at each pixel. Each filter of the MSFA samples the scene radiance in a given spectral band according to its spectral sensitivity function (SSF), so that each pixel of the acquired raw image represents a single band (see Fig. 1). The fully-defined MS image is estimated from the acquired raw values by demosaicing. To mitigate the impact of outdoor illumination variation on scene spectral signatures, the demosaiced image may undergo an additional post-processing step, such as white balancing or reflectance estimation [1].
@@ II Related works
II-A Learning from raw images
@@
If a descriptor effectively processes a raw image, it has the potential to achieve comparable or superior classification performances compared to demosaiced images because demosaicing generates artifacts that may compromise the quality of feature representation. Mihoubi et al. [28] proposed a method based on the local binary pattern (LBP) operator to compute texture features from raw images. The method analyzes the raw image with respect to the MSFA basic pattern and its band arrangement to build a texture descriptor by concatenating $B^{2}$ LBP histograms. Chen et al. [6] proposed histogram of oriented mosaic gradients (HOMG) to extract spatio-spectral features from raw images. First- and second-order horizontal and vertical gradient operators are designed to capture the spatial and spectral information. Then, the raw image is convolved with the designed gradient operators, resulting in spatio-spectral gradient maps. From these maps, four 31-dimensional histograms are constructed based on the magnitude and orientation of the gradients. The final 124-dimensional feature vector is obtained by concatenating these four histograms. Amziane et al. [2] proposed a CNN called MSFA-Net for feature extraction from raw reflectance images, guided by the MSFA basic pattern. Their model is composed of three different convolutional layers where the first layer captures spatio-spectral interactions between the channel values in each basic pattern, and two max-pooling layers in between. The produced feature maps are averaged channel-wise then fed into a fully-connected layer to provide the final 128-dimensional feature vector. In [12], a modified VGG-11 model pre-trained on ImageNet dataset is re-trained with raw images and used for raw texture feature extraction under different illumination conditions. Oppositely to [2], this approach applies a CNN directly to raw images without proper consideration of the basic pattern which may lead to potential inconsistency in spatio-spectral representation learning. In the same spirit of [2], we design a new model guided by the MSFA basic pattern to learn consistent and discriminant deep features from raw images.
III Method
@@
In this Section, we describe our method for illumination-robust learning from raw images. First, we describe our approach to white-balance (or correct) the raw radiance values to achieve raw spectral constancy. Then, we present our MSFA-preserving augmentations. Finally, we introduce RawMixer architecture for learning from the white-balanced raw images.
III-A Raw spectral constancy
@@
A single-sensor MS camera fitted with an MSFA provides a raw image $I^{\text{raw}}$ of size $X\times X$, where $X=m\cdot B$ is a multiple of the MSFA basic pattern width $B$. A single band $b\in\{0,...,B^{2}-1\}$ is associated with each pixel $p$ according to the MSFA. Under the Lambertian surface assumption and a global illumination source, $I^{\text{raw}}$ can be described as a function mainly dependent on three physical factors: the illumination relative spectral power distribution (RSPD) $L(\lambda)$, the surface element $s$ spectral reflectance $R_{s}(\lambda)$, and the sensor SSFs  $\{S^{b}(\lambda)\}_{b=0}^{B^{2}-1}$. By discarding the optical attenuation of the lens, the image integration time, and the quantization function, the radiance measured at pixel $p$ of the raw image $I^{\text{raw}}$ can be then simply expressed as:
@@
where the denominator is the estimated illumination associated to the $b$-th channel, and $\underset{5\times 5}{\text{median}}${} is a $5\times 5$ median filter to avoid corrupted and noisy pixels. Lastly, $\tilde{\textbf{O}}^{(B^{2})}$ is reorganized (shuffled) back to $\tilde{I}^{raw}$ using the shuffling operator $\mathcal{PS}:~{}{\mathbb{R}^{m\times m\times B^{2}}}\mapsto\mathbb{R}^{m\cdot B\times m\cdot B}$:
$\tilde{I}^{raw}_{p(x\cdot B,y\cdot B)}=\mathcal{PS}(\tilde{\textbf{O}}^{(B^{2})},x,y)=\tilde{O}^{\mathrm{MSFA}(p(x\cdot B,y\cdot B))}_{p(x,y)}\text{.}$ (6)
III-B MSFA-preserving augmentation
Traditional image augmentation techniques like rotation, cropping, and flipping are not directly applicable to raw images as they alter the structure of the basic pattern (see Fig. 2(a)). In [25], a method for Bayer pattern preservation during horizontal and vertical flips has been introduced. This involves column and/or row cropping after the augmentation to obtain the correct patterning of the raw image. However this approach is unsuitable for MSFAs with no redundant bands. To address this, we introduce seven augmentations tailored for MSFAs without redundant bands, applied to the unshuffled raw image instead of the original raw image. These transformations include MSFA-preserving vertical/horizontal flips, translations along x and y-axis, and texture remodeling. Additionally, we introduce Gaussian noise (directly to the raw image, with $\mu=0$ and $\sigma=0.25$), as a common artifact in raw image acquisition and optical distortion. Horizontal and vertical flips are performed by unshuffling the raw image (see Eq. (3)), applying the flip, and then shuffling it back (see Eq. (6)) to preserve the band arrangement. Translations are performed with step sizes that are multiples of the basic pattern width $B$ along the x and y-axis. Texture remodeling involves randomly replacing raw sub-patches matching the basic MSFA pattern with sub-patches from different locations within the raw image.
@@ III-B MSFA-preserving augmentation
III-C Raw feature extraction
We introduce RawMixer (see Fig. 3), an original hybrid network that combines the strengths of both transformers and CNNs to capture discriminant spatio-spectral interactions in raw images. The WB input raw image $\tilde{I}^{\mathrm{raw}}$ (see Eqs. (5) and (6)) of size $m\cdot B\times m\cdot B$ pixels is first fed to the raw ConvMixer $\Phi_{mixer}$ for spatio-spectral feature extraction. Its raw convolutional layer is designed to guide the feature extraction according to the MSFA basic pattern. It uses $320$ convolutional kernels $\{H_{n}\}_{n=0}^{319}$ of size $B\times B$ and depth 1, with a stride of $B$ pixels along both spatial dimensions and no padding. The $B$-pixel stride ensures that each kernel coefficient remains associated with the same MSFA band throughout all convolutions. This first layer learns spatio-spectral interactions among channel values within each raw basic patch that matches the basic MSFA pattern. The convolution between $\tilde{I}^{\mathrm{raw}}$ and a kernel $H_{n}$, $n\in[\![0,319]\!]$, is defined at each pixel $(x,y)\in[\![0,m-1]\!]^{2}$ as:
@@ III-C Raw feature extraction
The resulting feature maps $\{F_{n}\}_{n=0}^{319}$ of size $m\times m$ pixels are processed through the mixing block, employing depth-wise and point-wise convolutions to blend feature values across spatial locations and channels. Subsequently, the resulting feature maps are added to the previous ones through a residual connection before undergoing a $2\times 2$ maxpooling operation to yield a feature map of size $320\cdot(\lfloor\frac{m}{2}\rfloor\times\lfloor\frac{m}{2}\rfloor)$, that is reshaped and fed to the transformer encoder for attention computation. The resulting matrix is average pooled token-wise (each pixel is a token) to provide a 384-dimensional feature vector. This vector is then passed through a fully-connected layer to add non-linearity and reduce its dimensionality to obtain the final 128-dimensional texture feature vector, which serves as input to the softmax layer.
Figure 3: RawMixer architecture. $\Phi_{mixer}$ learns deep spatio-spectral interactions guided by the MSFA basic pattern. $\Phi_{t}$ is the (positional encoding free) transformer encoder that takes in the $320\cdot(\lfloor{\frac{m}{2}\rfloor}\times\lfloor{\frac{m}{2}\rfloor})$ feature maps provided by $\Phi_{mixer}$ reshaped as $\lfloor{\frac{m}{2}\rfloor}^{2}$ tokens $\times$ $320$ features. It learns another embedding through self-attention and feed-forward layers. SELU: scaled exponential linear unit, BN: batch normalization, FC: fully-connected layer. Filter depths (1 in raw conv. layer, 320 for ConvMixer layer) are not shown for sake of clarity.
IV Experiments
In this section, we investigate the effectiveness of our approach by conducting experiments on MS texture classification. For this purpose, we simulate raw radiance images based on the HyTexila database [21].
IV-A MSFAs
We consider three MSFAs (see Fig. 4) that are defined by repetition of a $B\times B$ basic pattern that samples $B^{2}$ different bands. We follow the MSFA arrangements in the VIS domain ($B=4$), NIR domain ($B=5$), and VIS-NIR [^1] IMEC VIS-NIR $2\times 2$ MSFA domain ($B=2$) of three snapshot cameras manufactured by IMEC [14].
@@ IV-B HyTexila database
Figure 5: RSPDs of the considered illuminants for all $\lambda\in[400\,\text{nm},1000\,\text{nm}]$.
IV-C Classification protocol
Patch extraction. Each image is horizontally split into two sub-images of size $508\times 1024$ ($B=2$ and $B=4$) or $510\times 1024$ ($B=5$) pixels. Non-overlapping square patches of width $X=n\cdot B$ are extracted from them, such as those associated to the second sub-image are used as test patches. We extract square patches of sizes $124^{2}$ and $64^{2}$ pixels (for $B=2$ and $B=4$), or $125^{2}$ and $65^{2}$ pixels ($B=5$). We perform MSFA-preserving augmentations (see Sec. III-B) to introduce some texture variations and to ensure enough patches to train the networks. We train on patches associated to D65 simulator and test on patches associated to extended A and Solar illuminants. To make the texture recognition more challenging, we perform horizontal/vertical flips, translation (with a step of $a\times B$ pixels, $a=\lfloor 0.6\cdot n\rfloor$) along y-axis, and texture remodeling augmentations only on learning patches, while the remaining ones (Gaussian noise, horizontal translation, and optical distortion) are performed on the test patches. The number $N_{l}$ of training patches is then $N_{l}\approx 27.4\cdot 10^{3}$ for $X\in\{64,65\}$, and $N_{l}\approx 8.9\cdot 10^{3}$ for $X\in\{124,125\}$, and the number $N_{t}$ of test patches is $N_{t}\approx 21.95\cdot 10^{3}$ for $X\in\{64,65\}$, and $N_{t}\approx 7.1\cdot 10^{3}$ for $X\in\{124,125\}$.

### v3

@@ Abstract
Multispectral (MS) snapshot cameras equipped with a MS filter array (MSFA), capture multiple spectral bands in a single shot, resulting in a raw mosaic image where each pixel holds only one channel value. The fully-defined MS image is estimated from the raw one through demosaicing, which inevitably introduces spatio-spectral artifacts. Moreover, training on fully-defined MS images can be computationally intensive, particularly with deep neural networks (DNNs), and may result in features lacking discrimination power due to suboptimal learning of spatio-spectral interactions. Furthermore, outdoor MS image acquisition occurs under varying lighting conditions, leading to illumination-dependent features. This paper presents an original approach to learn discriminant and illumination-robust features directly from raw images. It involves: raw spectral constancy to mitigate the impact of illumination, MSFA-preserving transformations suited for raw image augmentation to train DNNs on diverse raw textures, and raw-mixing to capture discriminant spatio-spectral interactions in raw images. Experiments on MS image classification show that our approach outperforms both handcrafted and recent deep learning-based methods, while also requiring significantly less computational effort. The source code is available at https://github.com/AnisAmziane/RawTexture.
{strip}[Uncaptioned image]
Figure 1: Fully-defined (a) vs. proposed (b) raw feature extraction pipeline. (a) requires estimating the fully-defined image first before post-processing and feature extraction, while (b) directly exploits the acquired raw image.
1 Introduction
Much of the recent success in several pattern recognition applications, such as crop/weed recognition [3, 34] and texture classification [29, 2], is thanks in part to the advancement of cameras and specifically to the emergence of compact "snapshot" MS cameras [17]. These cameras rely on an MSFA laid over the sensor to acquire several spectral bands in one shot and provide a raw mosaic image, in which a single channel value is available at each pixel. Each filter of the MSFA samples the scene radiance in a given spectral band according to its spectral sensitivity function (SSF), so that each pixel of the acquired raw image represents a single band (see Fig. 2). The fully-defined MS image is estimated from the acquired raw values by demosaicing. To mitigate the impact of outdoor illumination variation on scene spectral signatures, the demosaiced image may undergo an additional post-processing step, such as white balancing or reflectance estimation [1].
@@ 2 Related works
2.1 Learning from raw images
@@
If a descriptor effectively processes a raw image, it has the potential to achieve comparable or superior classification performances compared to demosaiced images because demosaicing generates artifacts that may compromise the quality of feature representation. Mihoubi et al. [29] proposed a method based on the local binary pattern (LBP) operator to compute texture features from raw images. The method analyzes the raw image with respect to the MSFA basic pattern and its band arrangement to build a texture descriptor by concatenating $B^{2}$ LBP histograms. Chen et al. [7] proposed histogram of oriented mosaic gradients (HOMG) to extract spatio-spectral features from raw images. First- and second-order horizontal and vertical gradient operators are designed to capture the spatial and spectral information. Then, the raw image is convolved with the designed gradient operators, resulting in spatio-spectral gradient maps. From these maps, four 31-dimensional histograms are constructed based on the magnitude and orientation of the gradients. The final 124-dimensional feature vector is obtained by concatenating these four histograms. Amziane et al. [2] proposed a CNN called MSFA-Net for feature extraction from raw reflectance images, guided by the MSFA basic pattern. Their model is composed of three different convolutional layers where the first layer captures spatio-spectral interactions between the channel values in each basic pattern, and two max-pooling layers in between. The produced feature maps are averaged channel-wise then fed into a fully-connected layer to provide the final 128-dimensional feature vector. In [13], a modified VGG-11 model pre-trained on ImageNet dataset is re-trained with raw images and used for raw texture feature extraction under different illumination conditions. Oppositely to [2], this approach applies a CNN directly to raw images without proper consideration of the basic pattern which may lead to potential inconsistency in spatio-spectral representation learning. Askary et al. [5] utilized the Places-CNN network, pre-trained on RGB images from the Places dataset. Three pseudo-RGB images are created from each raw image, and each is processed individually by a separate pre-trained Places-CNN. The final prediction is determined by combining the softmax probabilities from all three networks and selecting the class with the highest score. While this method exploits raw images, it does not directly learn from them. Moreover, it disrupts the original spatio-spectral scene structure present in each raw image. Furthermore, a separate network for each pseudo RGB image is required, leading to high computational demands and becomes impractical as the number of spectral bands grows (e.g., 8 networks for raw images obtained by a 25-band MSFA). In contrast to the aforementioned methods, our approach first mitigates the impact of illumination on raw images before being processed by the network. Additionally, we apply specific augmentations that maintain the MSFA basic pattern integrity throughout the augmentation process. Finally, we design a hybrid deep model guided by the MSFA basic pattern to learn consistent and discriminant deep features directly from raw images.
3 Method
@@
In this Section, we describe our method for illumination-robust learning from raw images. First, we describe our approach to white-balance (or correct) the raw radiance values to achieve raw spectral constancy. Then, we present our MSFA-preserving augmentations. Finally, we introduce RawMixer architecture for learning from the raw images.
3.1 Raw spectral constancy
@@
A single-sensor MS camera fitted with an MSFA provides a raw image $I^{\text{raw}}$ of size $X\times X$, where $X=m\cdot B$ is a multiple of the MSFA basic pattern width $B$, with $m$ an integer value representing the scaling factor for the image size relative to the basic pattern size $B$. A single band $b\in\{0,...,B^{2}-1\}$ is associated with each pixel $p$ according to the MSFA. Under the Lambertian surface assumption and a global illumination source, $I^{\text{raw}}$ can be described as a function mainly dependent on three physical factors: the illumination relative spectral power distribution (RSPD) $L(\lambda)$, the surface element $s$ spectral reflectance $R_{s}(\lambda)$, and the sensor SSFs  $\{S^{b}(\lambda)\}_{b=0}^{B^{2}-1}$. By discarding the optical attenuation of the lens, the image integration time, and the quantization function, the radiance measured at pixel $p$ of the raw image $I^{\text{raw}}$ can be then simply expressed as:
@@
where the denominator is the estimated illumination associated to the $b$-th channel, and $\underset{5\times 5}{\text{median}}${} is a $5\times 5$ median filter to avoid corrupted and noisy pixels. Lastly, $\tilde{\textbf{O}}^{(B^{2})}$ is reorganized (shuffled) back to $\tilde{I}^{\mathrm{raw}}$ using the shuffling operator $\mathcal{PS}:~{}{\mathbb{R}^{m\times m\times B^{2}}}\mapsto\mathbb{R}^{m\cdot B\times m\cdot B}$:
$\tilde{I}^{\mathrm{raw}}_{p(x\cdot B,y\cdot B)}=\mathcal{PS}(\tilde{\textbf{O}}^{(B^{2})},x,y)=\tilde{O}^{\mathrm{MSFA}(p(x\cdot B,y\cdot B))}_{p(x,y)}\text{.}$ (6)
3.2 MSFA-preserving augmentation
Traditional image augmentation techniques like rotation, cropping, and flipping are not directly applicable to raw images as they alter the structure of the basic pattern (see Fig. 3(a)). In [26], a method for Bayer pattern preservation during horizontal and vertical flips has been introduced. This involves column and/or row cropping after the augmentation to obtain the correct patterning of the raw image. However this approach is unsuitable for MSFAs with no redundant bands. To address this, we introduce seven augmentations tailored for MSFAs without redundant bands, applied to the unshuffled raw image instead of the original raw image. These transformations include MSFA-preserving vertical and horizontal flips, translation along x and y-axis, and texture remodeling. Additionally, we introduce Gaussian noise (directly to the raw image, with $\mu=0$ and $\sigma=0.25$), as a common artifact in raw image acquisition, and optical distortion. This distortion bends the image (barrel effects), causing straight lines to appear curved, much like the effect of a fish-eye lens. Horizontal and vertical flips are performed by unshuffling the raw image (see Eq. (3)), applying the flip, and then shuffling it back (see Eq. (6)) to preserve the band arrangement. Translations are performed with step sizes that are multiples of the basic pattern width $B$ along the x and y-axis. Texture remodeling involves randomly replacing raw sub-patches matching the basic MSFA pattern with sub-patches from different locations within the raw image.
@@ 3.2 MSFA-preserving augmentation
3.3 Raw feature extraction
We introduce RawMixer (see Fig. 4), a hybrid network that combines the strengths of both transformers and CNNs to capture discriminant spatio-spectral interactions in raw images. The WB input raw image $\tilde{I}^{\mathrm{raw}}$ (see Eqs. (5) and (6)) of size $m\cdot B\times m\cdot B$ pixels, is first fed to the raw ConvMixer $\Phi_{mixer}$ for spatio-spectral feature extraction. Its raw convolutional layer is designed to guide the feature extraction according to the MSFA basic pattern. It uses $320$ convolutional kernels $\{H_{n}\}_{n=0}^{319}$ of size $B\times B$ and depth 1, with a stride of $B$ pixels along both spatial dimensions and no padding. The $B$-pixel stride ensures that each kernel coefficient remains associated with the same MSFA band throughout all convolutions. This first layer learns spatio-spectral interactions among channel values within each raw basic patch that matches the basic MSFA pattern. The convolution between $\tilde{I}^{\mathrm{raw}}$ and a kernel $H_{n}$, $n\in[\![0,319]\!]$, is defined at each pixel $(x,y)\in[\![0,m-1]\!]^{2}$ as:
@@ 3.3 Raw feature extraction
The resulting feature maps $\{F_{n}\}_{n=0}^{319}$ of size $m\times m$ pixels are processed through the mixing block, employing depth-wise and point-wise convolutions to blend feature values across spatial locations and channels. Subsequently, the resulting feature maps are added to the previous ones through a residual connection before undergoing a $2\times 2$ maxpooling operation to yield a feature map of size $320\cdot(\lfloor\frac{m}{2}\rfloor\times\lfloor\frac{m}{2}\rfloor)$, that is reshaped and fed to the transformer encoder for attention computation. The resulting matrix is average pooled token-wise (each pixel is a token) to provide a 320-dimensional feature vector. This vector is then passed through a fully-connected layer to add non-linearity and reduce its dimensionality to obtain the final 128-dimensional texture feature vector, which serves as input to the softmax layer.
Figure 4: RawMixer architecture. $\Phi_{mixer}$ learns deep spatio-spectral interactions guided by the MSFA basic pattern. $\Phi_{t}$ is the (positional encoding free) transformer encoder that takes in the $320\cdot(\lfloor{\frac{m}{2}\rfloor}\times\lfloor{\frac{m}{2}\rfloor})$ feature maps provided by $\Phi_{mixer}$ reshaped as $\lfloor{\frac{m}{2}\rfloor}^{2}$ tokens $\times$ $320$ features. It learns another embedding through self-attention and feed-forward layers. The depth of the ConvMixer and transformer encoders is set to 2 in our experiments. SELU: scaled exponential linear unit, BN: batch normalization, FC: fully-connected layer. Filter depths (1 in raw conv. layer, 320 for ConvMixer layer) are not shown for sake of clarity.
4 Experiments
In this section, we investigate the effectiveness of our approach by conducting experiments on MS texture classification. For this purpose, we simulate raw radiance images based on the HyTexila [22] and SpecTex [30] datasets.
4.1 MSFAs
We consider three MSFAs (see Fig. 5) that are defined by repetition of a $B\times B$ basic pattern that samples $B^{2}$ different bands. We follow the MSFA arrangements in the VIS domain ($B=4$), NIR domain ($B=5$), and VIS-NIR [^1] IMEC VIS-NIR $2\times 2$ MSFA domain ($B=2$) of three snapshot cameras from IMEC [15].
@@ 4.2.1 HyTexila
Figure 6: RSPDs of the considered illuminants.
4.2.2 SpecTex
It contains 60 MS reflectance images of textures where each can be regarded as a single class. Each image has 77 channels in the range $[400\,\text{nm},780\,\text{nm}]$ of size $640\times 640$ pixels. We simulate raw radiance images following the same procedure used for HyTexila dataset.
4.3 Classification protocol
Patch extraction. For HyTexila, each image is horizontally split into two sub-images. To ensure that the first pixel (top-left corner) aligns with the first band of the MSFA, the sub-images are sized at either $508\times 1024$ (for $B=2$ and $B=4$) or $510\times 1024$ (for $B=5$) pixels. Non-overlapping square patches of width $X=n\cdot B$ are extracted from them, such as those associated to the second sub-image are used as test patches. We extract square patches of sizes $124^{2}$ and $64^{2}$ pixels (for $B=2$ and $B=4$), or $125^{2}$ and $65^{2}$ pixels ($B=5$). We perform MSFA-preserving augmentations (see Sec. 3.2) to introduce some texture variations and to ensure enough patches to train the networks. We train on patches associated to D65 simulator and test on patches associated to extended A and Solar illuminants. To make the texture recognition more challenging, we perform horizontal/vertical flips, translation (with a step of $a\times B$ pixels, $a=\lfloor 0.6\cdot n\rfloor$) along y-axis, and texture remodeling augmentations only on learning patches, while the remaining ones (Gaussian noise, horizontal translation, and optical distortion) are performed on the test patches. The number $N_{l}$ of training patches is then $N_{l}\approx 27.4\cdot 10^{3}$ for $X\in\{64,65\}$, and $N_{l}\approx 8.9\cdot 10^{3}$ for $X\in\{124,125\}$, and the number $N_{t}$ of test patches is $N_{t}\approx 21.95\cdot 10^{3}$ for $X\in\{64,65\}$, and $N_{t}\approx 7.1\cdot 10^{3}$ for $X\in\{124,125\}$.
For SpexTex, we also horizontally split each image into two sub-images of size $320\times 320$ pixels. Note that IMEC25 MSFA is not considered here because SpecTex images are defined only in the VIS domain, which corresponds to IMEC4 and IMEC16 spectral bands. Square patches of sizes $124^{2}$ and $64^{2}$ pixels (for $B=2$ and $B=4$) are extracted and augmented following the same procedure used for HyTexila. The number $N_{l}$ of training patches is then $N_{l}\approx 48\cdot 10^{2}$ for $X=64$, and $N_{l}\approx 38.4\cdot 10^{2}$ for $X=124$, and the number $N_{t}$ of test patches is $N_{t}\approx 21.95\cdot 10^{3}$ for $X=64$, and $N_{t}\approx 960$ for $X=124$.